OpenStudy (anonymous):

Divide. ((x^2-64)/(x^2)) / ((x^2-8x)/(x^2+2x-80)

OpenStudy (anonymous):

If we split this into terms.. 1: (x^2-64)/(x^2) 2: (x^2-8x)/(x^2+2x-80) To make this a bit less confusing, start by multiplying the numerator of the first by the denominator of the second and vice versa. In other words, flip one term and multiply top and bottom. So... (x^2)/(x^2-64) o (x^2-8x)/(x^2+2x-80)

OpenStudy (anonymous):

\[\frac{x^2-64}{x^2}\times \frac{x^2+2x-80}{x^2-8x}\] \[\frac{(x+8)(x-8)(x-8)(x+10)}{x^2\times x(x-8)}\] \[\frac{(x+8)(x-8)(x+10)}{x^3}\]